I just want to know did I get this right. I got 11?

Use order of operations to simplify.
4.2^2-20/4

Answer by richard1234(7193)

Do you mean 4.2^2-20/4 or is it problem 4, 2^2-20/4? I'm assuming the latter because I don't see how one could get 11 with a 4.2^2.

Without parentheses, 2^2-20/4 is interpreted as 4-20/4 = 4-5 = -1.